# On the Dirichlet problem for the degenerate $k$-Hessian equation

### v1

@@ 2. Preliminaries
$\frac{\left|h^{\prime}(0)\right|}{\sqrt{\beta}}\leq\sqrt{2}\sqrt{h^{\prime\prime}(0)}+2C\left(h^{\prime\prime}(0)\right)^{-\frac{3}{2}}\beta^{\frac{1}{2}}\sqrt{h^{\prime\prime}(0)}<\left(\sqrt{2}+2CK_{1}^{-\frac{3}{2}}\right)\sqrt{h^{\prime\prime}(0)}.$
Given that $\alpha<1/2$, choose $K_{1}$ large such that
@@
Some fundamental properties of $\sigma_{k}$ are listed below, and one can refer to [27, 14, 26, 33].
@@
For any $k\in\{1,2,\dots,n\}$ and $A\in\mathscr{S}^{n}$ with $\lambda(A)\in\Gamma_{k}$, the following hold for all $2\leq k\leq n$:
@@
$F^{ij}\big[D^{2}u\big]u_{eij}=\frac{p}{k}g^{\frac{p}{k}-\frac{1}{2}}\frac{\partial_{e}g}{\sqrt{g}}=\left(\frac{p}{k}g^{\frac{p}{k-1}-\frac{1}{2}}\frac{\partial_{e}g}{\sqrt{g}}\right)f^{\frac{-1}{k(k-1)}},\quad\forall e\in\mathbb{S}^{n-1}.$ (2.10)
A second differentiation combined with the concavity of $F$ yields
@@
$F^{ij}\big[D^{2}u\big]u_{eeij}\geq\frac{p}{k}g^{\frac{p}{k}-1}\left(\partial_{ee}g-\frac{k-p}{k}\frac{\left|\partial_{e}g\right|^{2}}{g}\right),\quad\forall e\in\mathbb{S}^{n-1}.$ (2.11)
@@
$\operatorname{tr}F^{ij}\big[D^{2}u\big]=\frac{1}{k}\sigma_{k}^{\frac{1-k}{k}}\big[D^{2}u\big]\sum_{i=1}^{n}\sigma_{k}^{ii}\big[D^{2}u\big]=\frac{n-k+1}{k}\sigma_{k}^{\frac{1-k}{k}}\big[D^{2}u\big]\sigma_{k-1}\big[D^{2}u\big].$
@@
$\left|\operatorname{tr}\left(GD^{2}u_{e}\right)\right|\leq C\sigma_{1}^{\frac{-1}{k-1}}\big[D^{2}u\big]g^{\frac{p}{k(k-1)}}\left|F^{ij}\big[D^{2}u\big]u_{eij}\right|,\quad\forall e\in\mathbb{S}^{n-1}.$
Then it follows from (2.10), (2.9), and (2.1) that
$\left|\operatorname{tr}\left(GD^{2}u_{e}\right)\right|\leq C\sigma_{1}^{\frac{-1}{k-1}}\big[D^{2}u\big]\frac{p}{k}g^{\frac{p}{k-1}-\frac{1}{2}}\frac{\left|\nabla g\right|}{\sqrt{g}}\leq K\sigma_{1}^{\frac{-1}{k-1}}\big[D^{2}u\big],\quad\forall e\in\mathbb{S}^{n-1},$
where $K$ depends on $\operatorname{dist}(\Omega,\partial\Omega_{0})$ and either $\|f^{1/(k-1)}\|_{C^{1,1}(\overline{\Omega_{0}})}$ or $\|f^{3/(2k-2)}\|_{C^{1,1}(\overline{\Omega_{0}})}$. Thus (2.15) is valid. ∎
@@
$\operatorname{tr}\left(GD^{2}u_{ee}\right)\geq-C\sigma_{1}^{\frac{-1}{k-1}}\big[D^{2}u\big]g^{\frac{p}{k(k-1)}}\frac{k-1}{k}g^{-\frac{1}{k}}\left|\partial_{ee}g-\frac{1}{k}\frac{\left|\partial_{e}g\right|^{2}}{g}\right|\geq-K_{1}\sigma_{1}^{\frac{-1}{k-1}}\big[D^{2}u\big]$
@@
3. $C^{1}$ estimate and the boundary tangential-normal estimate
Following Section 2 of Caffarelli-Nirenberg-Spruck [2], one can construct a subsolution $\underline{u}$ to equation (1.1) satisfying
$\begin{cases}\sigma_{k}\big(\lambda\big(D^{2}\underline{u}\big)\big)\geq f\quad\text{in}\ \Omega,\\ \underline{u}=\varphi\quad\text{on}\ \partial\Omega.\end{cases}$
Let $h$ be the harmonic function in $\Omega$ which equals $\varphi$ on $\partial\Omega$. Using the maximum principle, we find
$\underline{u}\leq u\leq h\quad\text{in}\ \Omega.$
Combining this with $u=\varphi$ on $\partial\Omega$, we obtain
@@ 3. $C^{1}$ estimate and the boundary tangential-normal estimate
where $C$ depends on $\Omega$, $\|f\|_{C^{0}(\overline{\Omega})}$, and $\|\varphi\|_{C^{3}(\partial\Omega)}$. Recall that $g=f^{3/(2k-2)}$ and p=(2k2)/3p=(2k-2)/3.
@@
$0\leq\sigma_{1}^{2}\big[D^{2}u\big]-\sum_{i,j=1}^{n}u_{ij}^{2}\leq\sigma_{1}^{2}\big[D^{2}u\big]-\sum_{j=1}^{n}u_{1j}^{2}=\sigma_{1}^{2}\big[D^{2}u\big]-4K^{2}|x|^{2}\leq\sigma_{1}^{2}\big[D^{2}u\big]-4K^{2}.$ (3.3)
@@
$F^{ij}\big[D^{2}u\big]u_{1ij}=\frac{p}{k}\frac{\partial_{1}g}{\sqrt{g}}f^{\frac{1}{4(k-1)}-\frac{1}{k(k-1)}}\geq-K_{1}f^{\frac{-1}{k(k-1)}}\quad\text{in}\ \Omega,$ (3.4)
where $K_{1}$ depends on $\operatorname{dist}(\Omega,\partial\Omega_{0})$ and $\|f^{3/(2k-2)}\|_{C^{1,1}(\overline{\Omega_{0}})}$.
@@
$0$ $\geq 2K\operatorname{tr}F^{ij}\big[D^{2}u\big]-K_{1}f^{\frac{-1}{k(k-1)}}$
$\geq\frac{2K}{C}\sigma_{1}^{\frac{1}{k-1}}\big[D^{2}u\big]f^{\frac{-1}{k(k-1)}}-K_{1}f^{\frac{-1}{k(k-1)}}$
@@
Since $u(x^{\prime},\rho(x^{\prime}))=\varphi(x^{\prime},\rho(x^{\prime}))$ on $\partial\Omega\cap\{|x^{\prime}|\leq r_{0}\}$, we derive
$0=(u-\varphi)_{\alpha}+(u-\varphi)_{n}\rho_{\alpha}=(u-\varphi)_{\alpha}+(u-\varphi)_{n}\sum_{\beta=1}^{n-1}\rho_{\alpha\beta}(0)x_{\beta}+O\Big(\big|x^{\prime}\big|^{2}\Big)$
on $\partial\Omega\cap\{|x^{\prime}|\leq r_{0}\}$ for all $1\leq\alpha\leq n-1$. Hence
$T_{\alpha}(u-\varphi)=-x_{n}\sum_{\beta=1}^{n-1}\rho_{\alpha\beta}(0)(u-\varphi)_{\beta}+O\Big(\big|x^{\prime}\big|^{2}\Big)=O\Big(\big|x^{\prime}\big|^{2}\Big)$ (3.6)
on $\partial\Omega\cap\{|x^{\prime}|\leq r_{0}\}$ for all $1\leq\alpha\leq n-1$. It follows from (2.10) and (2.1) that
$\left|F^{ij}[D^{2}u]\left(T_{\alpha}(u-\varphi)\right)_{ij}\right|\leq\left|T_{\alpha}(g^{p})\right|+C_{1}\operatorname{tr}F^{ij}\big[D^{2}u\big]\leq C_{2}f^{\frac{-1}{k(k-1)}}+C_{1}\operatorname{tr}F^{ij}\big[D^{2}u\big]$ (3.7)
@@
Define a function
@@
$F^{ij}\big[D^{2}u\big]\left(C_{3}(v_{ij}-\delta_{1}\delta_{ij})-u_{ij}\right)$ $\geq F\big[C_{3}\big(D^{2}v-\delta_{1}\mathbf{I}_{n}\big)\big]-F\big[D^{2}u\big]$
$=C_{3}\sigma_{k}^{\frac{1}{k}}\big[D^{2}v-\delta_{1}\mathbf{I}_{n}\big]-f^{\frac{1}{k}}$
$>0$
@@
$\left|\left(T_{\alpha}u\right)_{n}\right|\geq-Kv_{n}-\left|\left(T_{\alpha}\varphi\right)_{n}\right|=K\left(\beta_{1}-2x_{n}+2\rho\big(x^{\prime}\big)\right)-\left|\left(T_{\alpha}\varphi\right)_{n}\right|\geq\frac{\beta_{1}}{2}K-\left|\left(T_{\alpha}\varphi\right)_{n}\right|$
@@
Using (3.11), (3.7), (3.8), and (3.12), we obtain at the point $x_{0}$,
@@
Therefore, claim (3.9) holds with $K$ depending on $\Omega$, $\|\varphi\|_{C^{3}(\partial\Omega)}$, $\|u\|_{C^{1}(\overline{\Omega})}$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and $\|f^{3/(2k-2)}\|_{C^{1,1}(\overline{\Omega_{0}})}$. Since $\nabla_{x^{\prime}}u(0)=\nabla_{x^{\prime}}\varphi(0)=0$, we have
$T_{\alpha}(u-\varphi)(0)=0,\quad\forall 1\leq\alpha\leq n-1.$
From (3.9), it follows that
$\pm\left(T_{\alpha}(u-\varphi)\right)_{n}(0)\leq-Kv_{n}(0)=K\beta_{1}.$
The proof of Lemma 3.2 is complete. ∎
Combining (3.1), Lemmas 3.1 and 3.2, we establish Theorem 1.1.
4. Reduction of the second-order estimate to the boundary
In this section, inspired by Guan [9], Guan-Trudinger-Wang [12], and Jiao-Wang [17], we reduce the estimate of second-order derivatives to the boundary second-order normal estimate, that is Theorem 1.2. Recall that $g=f^{3/(2k-2)}$ and p=(2k2)/3p=(2k-2)/3.
Proof of Theorem 1.2.
@@ 4. Reduction of the second-order estimate to the boundary
$0\geq C(\Delta u)^{\frac{1}{k-1}}f^{1-\frac{1}{k-1}}+pg^{p-1}\sum_{l=1}^{n}\left[(p-1)\frac{|\partial_{x_{l}}g|^{2}}{g}+\partial_{x_{l}x_{l}}g\right]-\sum_{l=1}^{n}\sigma_{k}^{ij,st}\big[D^{2}u\big]u_{ijl}u_{stl}.$ (4.3)
@@
$\frac{\sigma_{k}^{ij,st}u_{ije}u_{ste}}{\sigma_{k}}\leq\frac{k-2}{k-1}\left(\frac{\sigma_{k}^{ij}u_{ije}}{\sigma_{k}}\right)^{2}+\frac{2}{k-1}\frac{\sigma_{k}^{ij}u_{ije}}{\sigma_{k}}\frac{\sigma_{1}^{ij}u_{ije}}{\sigma_{1}}-\frac{k}{k-1}\left(\frac{\sigma_{1}^{ij}u_{ije}}{\sigma_{1}}\right)^{2}$

### v2

@@ 2. Preliminaries
$\frac{\left|h^{\prime}(0)\right|}{\sqrt{\beta}}$ $\leq\sqrt{2}\sqrt{h^{\prime\prime}(0)}+2C\left(h^{\prime\prime}(0)\right)^{-\frac{3}{2}}\beta^{\frac{1}{2}}\sqrt{h^{\prime\prime}(0)}$
$<\left(\sqrt{2}+2CK_{1}^{-\frac{3}{2}}\right)\sqrt{h^{\prime\prime}(0)}.$
Given that $\alpha<1/2$, choose $K_{1}$ sufficiently large such that
@@
Remark 2.3.
In general, it is impossible to improve the exponent $\alpha$ to any $\alpha>1/2$. Consider the family of functions $g_{\beta}(x):=(x+\beta^{1/2})^{2}$, where $\beta>0$ is a parameter. It is clear that $g_{\beta}>0$ and $g_{\beta}\in C^{\infty}(\mathbb{R})$. Computing (2.2) at $x=0$ and letting $\beta\rightarrow 0$ yield that $\alpha\leq 1/2$.
By convention, we set $\sigma_{0}(\lambda)\equiv 1$. Some fundamental properties of $\sigma_{k}$ are listed below, and one can refer to [27, 14, 26, 33].
@@
For any $k\in\{2,3,\dots,n\}$ and any symmetric matrix $A$ with $\lambda(A)\in\Gamma_{k}$, the following hold:
@@
$F^{ij}\big[D^{2}u\big]u_{eij}=\frac{p}{k}g^{\frac{p}{k}-\frac{1}{2}}\frac{\partial_{e}g}{\sqrt{g}}=\left(\frac{p}{k}g^{\frac{p}{k-1}-\frac{1}{2}}\frac{\partial_{e}g}{\sqrt{g}}\right)f^{\frac{-1}{k(k-1)}}$ (2.10)
for all $e\in\mathbb{S}^{n-1}$. A second differentiation combined with the concavity of $F$ yields
@@
$F^{ij}\big[D^{2}u\big]u_{eeij}\geq\frac{p}{k}g^{\frac{p}{k}-1}\left(\partial_{ee}g-\frac{k-p}{k}\frac{\left|\partial_{e}g\right|^{2}}{g}\right)$ (2.11)
for all $e\in\mathbb{S}^{n-1}$.
@@
$\operatorname{tr}F^{ij}\big[D^{2}u\big]$ $=\frac{1}{k}\sigma_{k}^{\frac{1-k}{k}}\big[D^{2}u\big]\sum_{i=1}^{n}\sigma_{k}^{ii}\big[D^{2}u\big]$
$=\frac{n-k+1}{k}\sigma_{k}^{\frac{1-k}{k}}\big[D^{2}u\big]\sigma_{k-1}\big[D^{2}u\big].$
@@
$\left|\operatorname{tr}\left(GD^{2}u_{e}\right)\right|\leq C\sigma_{1}^{\frac{-1}{k-1}}\big[D^{2}u\big]g^{\frac{p}{k(k-1)}}\left|F^{ij}\big[D^{2}u\big]u_{eij}\right|$
for all $e\in\mathbb{S}^{n-1}$. Then it follows from (2.10), (2.9), and (2.1) that
$\left|\operatorname{tr}\left(GD^{2}u_{e}\right)\right|\leq C\sigma_{1}^{\frac{-1}{k-1}}\big[D^{2}u\big]\frac{p}{k}g^{\frac{p}{k-1}-\frac{1}{2}}\frac{\left|\nabla g\right|}{\sqrt{g}}\leq K\sigma_{1}^{\frac{-1}{k-1}}\big[D^{2}u\big]$
for all $e\in\mathbb{S}^{n-1}$, where the constant $K$ depends on $\operatorname{dist}(\Omega,\partial\Omega_{0})$ and either $\|f^{1/(k-1)}\|_{C^{1,1}(\overline{\Omega_{0}})}$ or $\|f^{3/(2k-2)}\|_{C^{1,1}(\overline{\Omega_{0}})}$. Thus (2.15) is valid. ∎
@@
$\operatorname{tr}\left(GD^{2}u_{ee}\right)$ $\geq-C\sigma_{1}^{\frac{-1}{k-1}}\big[D^{2}u\big]g^{\frac{p}{k(k-1)}}\frac{k-1}{k}g^{-\frac{1}{k}}\left|\partial_{ee}g-\frac{1}{k}\frac{\left|\partial_{e}g\right|^{2}}{g}\right|$
$\geq-K_{1}\sigma_{1}^{\frac{-1}{k-1}}\big[D^{2}u\big]$
@@
3. Reduction to the boundary second-order normal estimate
It follows from Caffarelli-Nirenberg-Spruck [2] that
@@ 3. Reduction to the boundary second-order normal estimate
where $C$ depends on $\Omega$, $\|f\|_{C^{0}(\overline{\Omega})}$, and $\|\varphi\|_{C^{3}(\partial\Omega)}$, but is independent of $\inf_{\Omega_{0}}f$.
@@
$0\leq\sigma_{1}^{2}\big[D^{2}u\big]-\sum_{j=1}^{n}u_{1j}^{2}\leq\sigma_{1}^{2}\big[D^{2}u\big]-4K^{2}.$ (3.3)
@@
$F^{ij}\big[D^{2}u\big]u_{1ij}=\frac{p}{k}\frac{\partial_{1}g}{\sqrt{g}}f^{\frac{1}{4(k-1)}-\frac{1}{k(k-1)}}\geq-K_{1}f^{\frac{-1}{k(k-1)}}$ (3.4)
in $\Omega$, where $K_{1}$ depends on $\operatorname{dist}(\Omega,\partial\Omega_{0})$ and $\|f^{3/(2k-2)}\|_{C^{1,1}(\overline{\Omega_{0}})}$.
@@
$0$ $\geq\frac{2K}{C}\sigma_{1}^{\frac{1}{k-1}}\big[D^{2}u\big]f^{\frac{-1}{k(k-1)}}-K_{1}f^{\frac{-1}{k(k-1)}}$
@@
It is standard that
$T_{\alpha}(u-\varphi)=O\Big(\big|x^{\prime}\big|^{2}\Big),\quad\forall 1\leq\alpha\leq n-1,$ (3.6)
on $\partial\Omega\cap\{|x^{\prime}|\leq r_{0}\}$. It follows from (2.10) and (2.1) that
$\left|F^{ij}[D^{2}u]\left(T_{\alpha}(u-\varphi)\right)_{ij}\right|$ $\leq\left|T_{\alpha}(g^{p})\right|+C_{1}\operatorname{tr}F^{ij}\big[D^{2}u\big]$
$\leq C_{2}f^{\frac{-1}{k(k-1)}}+C_{1}\operatorname{tr}F^{ij}\big[D^{2}u\big]$ (3.7)
@@
$\omega_{r}:=\left\{x\in\Omega:\ \rho\big(x^{\prime}\big)<x_{n}<\rho\big(x^{\prime}\big)+r^{4},\ \big|x^{\prime}\big|<r\right\}$
for all $r\leq r_{0}$. Define a function
@@
$F^{ij}\big[D^{2}u\big]\left(C_{3}(v_{ij}-\delta_{1}\delta_{ij})-u_{ij}\right)\geq C_{3}\sigma_{k}^{\frac{1}{k}}\big[D^{2}v-\delta_{1}\mathbf{I}_{n}\big]-f^{\frac{1}{k}}>0$
@@
$\left|\left(T_{\alpha}u\right)_{n}\right|\geq-Kv_{n}-\left|\left(T_{\alpha}\varphi\right)_{n}\right|$ $=K\left(\beta_{1}-2x_{n}+2\rho\big(x^{\prime}\big)\right)-\left|\left(T_{\alpha}\varphi\right)_{n}\right|$
$\geq\frac{\beta_{1}}{2}K-\left|\left(T_{\alpha}\varphi\right)_{n}\right|$
@@
Using (3.11), (3), (3.8), and (3.12), we obtain at the point $x_{0}$,
@@
Therefore, claim (3.9) holds with $K$ depending on $\Omega$, $\|\varphi\|_{C^{3}(\partial\Omega)}$, $\|u\|_{C^{1}(\overline{\Omega})}$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and $\|f^{3/(2k-2)}\|_{C^{1,1}(\overline{\Omega_{0}})}$. It follows immediately that
$\pm\left(T_{\alpha}(u-\varphi)\right)_{n}(0)\leq-Kv_{n}(0)=K\beta_{1}$
for all $1\leq\alpha\leq n-1$. The proof of Lemma 3.2 is complete. ∎
Inspired by Guan [9], Guan-Trudinger-Wang [12], and Jiao-Wang [17], we reduce the estimate of second-order derivatives to the boundary.
Lemma 3.3.
Proof.
@@
$0$ $\geq C(\Delta u)^{\frac{1}{k-1}}f^{1-\frac{1}{k-1}}+pg^{p-1}\sum_{l=1}^{n}\left[(p-1)\frac{|\partial_{x_{l}}g|^{2}}{g}+\partial_{x_{l}x_{l}}g\right]$ (3.15)
$\hskip 12.80373pt-\sum_{l=1}^{n}\sigma_{k}^{ij,st}\big[D^{2}u\big]u_{ijl}u_{stl}.$
@@
$\frac{\sigma_{k}^{ij,st}u_{ije}u_{ste}}{\sigma_{k}}$ $\leq\frac{k-2}{k-1}\left(\frac{\sigma_{k}^{ij}u_{ije}}{\sigma_{k}}\right)^{2}$
$\hskip 12.80373pt+\frac{2}{k-1}\frac{\sigma_{k}^{ij}u_{ije}}{\sigma_{k}}\frac{\sigma_{1}^{ij}u_{ije}}{\sigma_{1}}-\frac{k}{k-1}\left(\frac{\sigma_{1}^{ij}u_{ije}}{\sigma_{1}}\right)^{2}$